# Automated Cell Structure Extraction for 3D Electron Microscopy by Deep Learning

### v1

@@ 3.1 Segmentation Anything Model
We used an ImageJ plugin to process the continuous cross-sectional images captured by FIB-SEM and applied the 3D Watershed algorithm. This technique enabled us to identify different cells individually, as shown in Figure 5-f. Thus, the process of creating an SBF image containing a single cell and identifying each cell and cell organelle, previously done manually by experts, has now been fully automated.
4 Discussion
@@
We have successfully established a method to automatically segment and focus on specific cells within a mixed spatial environment in 3D imagery. By incorporating the latest deep learning models, which are being released at an ever-increasing rate, we can further enhance the accuracy of our inferences. This research method was achieved by employing deep learning analysis techniques, which require a substantial amount of high-quality, expertly labeled data as a prerequisite. The necessity for extensive labeled data poses a potential bottleneck for future advancements in this field. This research holds significant promise for biologists, particularly in aiding the morphological and dynamic analysis of cellular and tissue structures in 3D. The ability to automatically segment and analyze intricate cellular compositions in three dimensions provides a powerful tool for biological research.
Tables
@@
Table 1: The dataset used for cross-validation of pre-training. ’Case Number’ refers to the consecutive number for each case of cross-validation. The number of cells used for training data in each cell cycle is denoted as ’Cells,’ while the number of cross-sectional images used for training is indicated as Learning Images ’LI.’ Validation Images ’VI.’ represent the number of cross-sectional images used for the validation data.
@@
Table 2: The scanning 3D U-Net Architecture (K, O, I, D, W, H are kernel size, output channel, input channel, output depth, output width, and output height, respectively).
@@
Table 3: Inference accuracy of the pre-trained model. Evaluated the inference results of the model for each cellular organelle, in each case of cross-validation and for each stage of the cell cycle, based on four criteria: IoU score.

### v3

@@ Results
We used an ImageJ plugin to process the continuous cross-sectional images captured by the FIB-SEM and applied the 3D watershed algorithm. This technique enabled us to identify different cells individually, as shown in Figure 6 (f). Thus, the process of creating an SBF image containing a single cell and identifying each cell and cell organelle, previously done manually by experts, was fully automated.
Discussion
@@
We successfully established a method of automatically segmenting and focusing on specific cells within a mixed spatial environment in 3D imagery. By incorporating the latest deep learning models, which are being released at an ever-increasing rate, we can further enhance the accuracy of our inferences. This research method was achieved by employing deep learning analysis techniques, which require a substantial amount of high-quality, expertly labeled data as a prerequisite. The necessity for extensive labeled data poses a potential bottleneck for future advancements in this field. This research holds significant promise for biologists, particularly in aiding the morphological and dynamic analysis of cellular and tissue structures in 3D. The ability to segment and analyze intricate cellular compositions automatically in three dimensions provides a powerful tool for biological research.
In principle, the technology presented here could be applied to images obtained by different types of 3D imaging devices, not limited to FIB-SEM. However, it appears especially effective to use machine learning for processing electron microscope images that contain only brightness information. A large volume of high-quality annotated data corresponding to 3D images is required to utilize this technology. This poses the most significant challenge when applying the technology to other cell types. Recently, various methods have been proposed for creating label data (semi-)manually and efficiently, and for performing zero-shot segmentation without label data. By integrating these methods, we may further enhance the technology described here to extract only specific structures from the environment crowded by multiple cells. In addition, it was suggested that label data of diverse patterns of division would be required when performing structure extraction for various cell states such as cell division, where the determination of the cell state would be another target of machine learning.
Data availability
The analysis codes and the raw data of the analysis results in this study are available from the corresponding author (Y. T.) on reasonable request.
Acknowledgements
This work was supported by JSPS KAKENHI grant number JP21K03487.
Author contributions statement
J. K. conceived and implemented the analysis algorithms. A. H. I. provided the 3D cell images. Y. T. supervised the research. All authors wrote and reviewed the manuscript.
Additional information
The authors declare no competing interests.
@@ Acknowledgements
Table 1: Dataset used for cross-validation of training. “Case Number” refers to the consecutive number for each case of cross-validation. The number of cells used for training data in each cell cycle is denoted as “Cells”, and the number of cross-sectional images used for training is indicated as “LI”. “VI” represents the number of cross-sectional images used for the validation data.
@@
Table 2: Scanning 3D U-Net architecture (K, O, I, D, W, and H are the kernel size, output channel, input channel, output depth, output width, and output height, respectively).
@@
Table 3: Inference accuracy of the trained model. Evaluation of the inference results of the model for each cellular organelle, in each case of cross-validation and for each stage of the cell cycle, based on the IoU score (1).
@@
Figure 1: Procedure to reconstruct a 3D model of a cell from sequential cross-sectional images. (a) SBF image that centrally includes the cell of interest, cropped from FIB-SEM images. (b) By reconstructing the data and removing the background, a 3D model of the cell is produced. (c) Three-dimensional organelle model visualized by further removing the cytoplasm, indicating cytoplasm (white), plastid (green), mitochondrion (yellow), peroxisome (orange), and nucleus (gray). (d) This SBF consists of sequential cross-sectional images, as depicted in the top row. Various organelles, the cytoplasm, and the background in these images were manually identified by experts, as shown in the bottom row. In this panel, images at six different depths (in the direction perpendicular to the images) are presented.
Figure 2: Structural changes associated with the cell cycle of C. merolae. Each structure includes the cytoplasm (white), plastid (green), mitochondrion (yellow), peroxisome (orange), and nucleus (gray). Each type of organelle is equally distributed to the daughter cells during division [6, 7]. The division of organelle structures occurs at nearly consistent timings regardless of the individual, although the timing varies depending on the type of organelle [13].
Figure 4: Examples of inference results for the validation dataset using a model obtained through the training. Each structure includes the cytoplasm (white), plastid (green), mitochondrion (yellow), peroxisome (orange), and nucleus (gray). The output images of the plastid and cytoplasm closely match the ground truth.
Figure 5: Cross-sections and label data of cells in various directions. (a) Three-dimensional cell model reconstructed from captured consecutive cross-sectional images. This model is visualized from the captured direction. (b) Consecutive cross-sectional images taken near the center of the cell (left) and label data created by experts based on them (right). (c) and (d) New cross-sectional images generated from stacked consecutive cross-sectional images (left) and the corresponding label data (right). These images were generated from two directions, one aligned with the captured direction and the other perpendicular. Each panel includes the cytoplasm (white), plastid (green), mitochondrion (yellow), peroxisome (orange), and nucleus (gray).
Figure 6: Schematic diagram of automatic identification of individual cells from a 3D image containing multiple cells using SAM and the 3D watershed method. (a), (b) Cross-sectional image of a sample containing multiple cells and its SBF. (c) Extraction of cells from the background by SAM. (d) Classification of cellular organelles by SAM. (e) Removal of the background identified by SAM from the SBF in (b). (f) Identification and color-coding of individual cells by the 3D watershed method.